# FPGA-Enabled Machine Learning Applications in Earth Observation: A Systematic Review

### v1

@@ Abstract.
New UAV technologies and the NewSpace era are transforming Earth Observation missions and data acquisition. Numerous small platforms generate large data volume, straining bandwidth and requiring onboard decision-making to transmit high-quality information in time. While Machine Learning allows real-time autonomous processing, FPGAs balance performance with adaptability to mission-specific requirements, enabling onboard deployment. This review systematically analyzes 66 experiments deploying ML models on FPGAs for Remote Sensing applications. We introduce two distinct taxonomies to capture both efficient model architectures and FPGA implementation strategies. For transparency and reproducibility, we follow PRISMA 2020 guidelines and share all data and code at https://github.com/CedricLeon/Survey_RS-ML-FPGA.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Computing methodologies Machine learningccs: Hardware Reconfigurable logic and FPGAsccs: Applied computing Earth and atmospheric sciences
@@ 1. Introduction
The last decade of research has seen major innovations in Machine Learning (ML), particularly in Deep Learning (DL). Every discipline has seen publications leveraging ML’s abstract learning capabilities to alleviate specific problems of the field. Nowadays, many state-of-the-art solutions incorporate ML methods in their pipeline. For instance, automatic navigation, for cars and Unmanned Aerial Vehicles (UAVs) alike, heavily relies on scene segmentation and obstacle detection. Similarly, video and image encoding apply learnable heuristics to select optimal patterns and splits in the data. Beyond vision, language processing benefits from the wide context available through transformer tokenization, significantly enhancing performance. By and large, these AI technologies have significant computational costs, making them inadequate in resource-constrained environments. Therefore, many applications deploying their algorithms close to the sensor, e.g., cars and UAVs On-Board Computers (OBCs), GoPros, or intelligent microphones, benefit from lightweight optimizations of these computationally intensive models. Approximate computing optimizations, trading computation load over accuracy, such as quantization, can typically reduce memory overhead by a factor of $4$ and computational costs by a factor of $16$ (Nagel et al., 2021).
In this survey, we take a systematic look at one application of ML for computer vision: Remote Sensing (RS) for Earth Observation (EO). Contemporary challenges, such as the climate crisis, urban planning, or defense concerns, transformed RS into an essential field. In particular, the rising NewSpace [^1] https://philab.esa.int/tag/new-space/ industry has decreased the costs and the production time of SmallSats. Because deploying CubeSats has never been so easy, we can observe a steady increase in the number of missions—to date, over 2600 nanosatellites have been launched (Kim et al., 2024). However, payload restrictions on these compact satellites reduce their onboard compute capacity relative to their larger counterparts, forcing them to run application-specific routines in flight and underscoring the need for efficient onboard data processing.
@@
In the particular example of RS, while the quantity of data grows, the bandwidth resources for the data transmission (downlink) stagnate (Pitonak et al., 2022). Because of such limitations, most current missions acquire data on-demand or over certain pre-configured conditions, inevitably missing out relevant information. The increasing sensing resolution and number of in-orbit satellites, paired with further restrictions on the radio-frequency spectrum, will worsen the problem (Růžička et al., 2022). Processing the data directly onboard, transmitting only compressed or application-tailored information, constitutes an elegant solution for future missions. Equipping future missions with edge computing platforms opens up possibilities to relieve data processing stress. Because SmallSats’ standard payloads typically include FPGA-based hardware (Siegle et al., 2015; George and Wilson, 2018), this survey focuses on analyzing the literature deploying ML-based solutions for RS applications on FPGAs.
@@
  • Systematic literature collection and analysis ensuring a rigorous and transparent review process following the PRISMA 2020 guidelines.
  • Two specialized taxonomies, one structuring RS applications based on their ML problem formulation, and another categorizing FPGA implementations by frameworks and design patterns.
  • Comprehensive evaluation of existing works, structured around eight Research Questions (RQs) to analyze the research landscape and method synergies.
  • Identification of research gaps, structured under five subtopics to guide future advancements in FPGA-enabled ML for Earth Observation.
  • Guidelines and recommendations to facilitate benchmarking, enhance reproducibility, and drive future research efforts.
@@ 1.1. Background
Remote Sensing (RS) [^2] Throughout this survey we will prefer the term Remote Sensing (RS) rather than Earth Observation (EO). RS specifically refers to acquiring information about the Earth’s surface from a distance while EO includes both remote sensing and in situ measurements. is the science of obtaining reliable information about objects or areas on the Earth’s surface without direct physical contact (Campbell and Wynne, 2011). It involves the acquisition, processing, and interpretation of data captured by sensors that detect electromagnetic radiation. The discipline is broadly divided into active and passive approaches. Active systems, such as radar and LiDAR, emit their own energy signals and record the returning echoes, while passive systems depend on external sources of energy, primarily sunlight. Furthermore, RS platforms are classified by their operational altitude into spaceborne and airborne systems. Spaceborne platforms, typically mounted on satellites, provide extensive temporal and spatial coverage. They enable continuous monitoring over vast areas and across diverse time scales—from near real-time updates to the analysis of long-term environmental trends (Zhu et al., 2017). Airborne platforms, on the other hand, offer higher spatial resolution over more localized areas, making them ideal for detailed studies that require fine-grained spatial information.
@@ 2.1. Research Questions
Finally in Section 5, R7-8 take a holistic view of the accomplishments in this field, discussing key achievements and, in line with the current NewSpace era, the potential of AI on the edge for EO missions.
  • RQ7: What combination of Machine Learning model, optimization strategies, and FPGA design best synergizes to optimize key Remote Sensing metrics?
  • RQ8: How are AI-powered edge computing solutions transforming EO missions?
To answer these research questions, we conduct a systematic search procedure adapted from the method proposed by the PRISMA 2020 guidelines (Page et al., 2021). The procedure starts by defining a prompt to frame for the population of records of interest. After collecting the results from the chosen research databases, we perform an initial screening pass via metadata analysis to filter out undesired records. The second screening pass involves reviewing the complete records, excluding any not clearly identified as being outside the survey’s scope based on the metadata alone.
@@ 2.2. Search Procedure
TOPICS: [ (Machine Learning OR Deep Learning OR Neural Network OR Convolutional Neural Network OR Recurrent Neural Network OR ML OR DL OR NN OR CNN OR RNN)
AND
@@
OR Remote Sensing OR Earth Science OR LiDAR OR SAR OR UAV OR Sentinel)
AND
@@
YEAR PUBLISHED: $2014$-$2024$
LANGUAGE: ENGLISH
The central part of the prompt consists of three expressions framing for any ML model, the usage of RS data in EO, and the explicit mention of FPGA devices. We choose not to consider articles published more than 10 years ago. Indeed, the FPGA technology, the state-of-the-art DL models, and even the tackled problems have evolved rapidly. Therefore, work from more than 10 years ago is often too different from modern studies to be compared fairly.
After retrieving the $119$ results of the prompt, we proceed to screen the articles to filter mismatches. To this effect, we pick 4 Exclusion Criteria (EC):
@@
While EC1 and EC2 are self-explanatory, EC3 mostly rules out non-EO applications of remote sensing technology, e.g., LiDAR for autonomous driving or SAR calibration. EC4 excludes studies with no experiments or details about the FPGA implementation. For example, EC4 eliminates records mentioning FPGA hardware as a possible solution or using a different FPGA acronym, e.g., fast patch-free global learning. We design this filtering process to retain only the most relevant, experiment-based studies that reflect the current state of FPGA-based ML in Remote Sensing.
After screening, $46$ articles are kept for review. Each article included in the survey is thoroughly read and tagged in Zotero (Corporation for Digital Scholarship, 2024). In particular, we use Zotero’s tagging system and API to systematically report the experiments’ specificities and performance metrics. The entire database extracted via Zotero’s API, along with the code used for analysis, figures, and main taxonomy tables, is available on GitHub at https://github.com/CedricLeon/Survey_RS-ML-FPGA. Our repository also hosts supplemental material, such as extended methodological considerations.
@@ 3. Mapping the Research Landscape: Applications, Models, and Hardware
In this section, we analyze the motivations, tackled problems, and solutions of each study included in the survey. In particular, we propose a taxonomy based on the ML and RS components of the experiments in Table 2.

### v2

@@ Abstract.
New UAV technologies and the NewSpace era are transforming Earth Observation missions and data acquisition. Numerous small platforms generate large data volume, straining bandwidth and requiring onboard decision-making to transmit high-quality information in time. While Machine Learning allows real-time autonomous processing, FPGAs balance performance with adaptability to mission-specific requirements, enabling onboard deployment. This review systematically analyzes $68$ experiments deploying ML models on FPGAs for Remote Sensing applications. We introduce two distinct taxonomies to capture both efficient model architectures and FPGA implementation strategies. For transparency and reproducibility, we follow PRISMA 2020 guidelines and share all data and code at https://github.com/CedricLeon/Survey_RS-ML-FPGA.
@@
copyright: ccjournal: CSURjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3800686isbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Computing methodologies Machine learningccs: Hardware Reconfigurable logic and FPGAsccs: Applied computing Earth and atmospheric sciences
Figure 1. Scope and content of the survey.
@@ 1. Introduction
The last decade of research has seen major innovations in Machine Learning (ML), particularly in Deep Learning (DL). Every discipline has seen publications leveraging ML’s abstract learning capabilities to alleviate specific problems of the field. Nowadays, many state-of-the-art solutions incorporate ML methods in their pipeline. For instance, automatic navigation, for cars and Unmanned Aerial Vehicles (UAVs) alike, heavily relies on scene segmentation and obstacle detection. Similarly, video and image encoding apply learnable heuristics to select optimal patterns and splits in the data. Beyond vision, language processing benefits from the wide context available through transformer tokenization, significantly enhancing performance. By and large, these AI technologies have significant computational costs, making them inadequate in resource-constrained environments. Therefore, many applications deploying their algorithms close to the sensor, e.g., cars and UAVs On-Board Computers (OBCs), GoPros, or intelligent microphones, benefit from lightweight optimizations of these computationally intensive models.
@@
In this survey, we take a systematic look at one application of ML for computer vision: Remote Sensing (RS) for Earth Observation (EO) [^1] Throughout this survey, we will prefer the term RS, as it is the core technology of EO, which also encompasses in-situ measurements. . Contemporary challenges, such as the climate crisis, urban planning, or defense concerns, transformed RS into an essential field. In particular, the rising NewSpace [^2] https://philab.esa.int/tag/new-space/ industry has decreased the costs and the production time of SmallSats. Because deploying CubeSats has never been so easy, we can observe a steady increase in the number of missions—to date, over 2600 nanosatellites have been launched (Kim et al., 2024). However, payload restrictions on these compact satellites reduce their onboard compute capacity relative to their larger counterparts, forcing them to run application-specific routines in flight and underscoring the need for efficient onboard data processing. In the particular example of RS, while the quantity of data grows, the bandwidth resources for the data transmission (downlink) stagnate (Pitonak et al., 2022). Because of such limitations, most current missions acquire data on-demand or over certain pre-configured conditions, inevitably missing out relevant information. The increasing sensing resolution and number of in-orbit satellites, paired with further restrictions on the radio-frequency spectrum, will worsen the problem (Růžička et al., 2022). Processing the data directly onboard, transmitting only compressed or application-tailored information, constitutes an elegant solution for future missions. Equipping future missions with edge computing platforms opens up possibilities to relieve data processing stress. Because SmallSats’ standard payloads typically include FPGA-based hardware (Siegle et al., 2015; George and Wilson, 2018), this survey focuses on analyzing the literature deploying ML-based solutions for RS applications on FPGAs (Fig. 1).
@@
  • Systematic literature collection and analysis ensuring a rigorous and transparent review process following the PRISMA 2020 guidelines (Section 2).
  • Comprehensive evaluation of existing works, structured around eight Research Questions (RQs) to analyze the research landscape (Section 3), the design strategies (Section 4), and method synergies (Section 5).
  • Two specialized taxonomies, one structuring RS applications based on their ML problem formulation (Section 3), and another categorizing FPGA implementations by frameworks and design patterns (Section 4).
  • Identification of research gaps, structured under five subtopics to guide future advancements in FPGA-enabled ML for Earth Observation (Section 6).
  • Guidelines and recommendations to facilitate benchmarking, enhance reproducibility, and drive future research efforts (Section 6).
@@ 1.1. Background
Remote Sensing (RS) is the science of obtaining reliable information about objects or areas on the Earth’s surface without direct physical contact (Campbell and Wynne, 2011). It involves the acquisition, processing, and interpretation of data captured by sensors that detect electromagnetic radiation. The discipline is broadly divided into active and passive approaches. Active systems, such as radar and LiDAR, emit their own energy signals and record the returning echoes, while passive systems depend on external sources of energy, primarily sunlight. Furthermore, RS platforms are classified by their operational altitude into spaceborne and airborne systems. Spaceborne platforms, typically mounted on satellites, provide extensive temporal and spatial coverage. They enable continuous monitoring over vast areas and across diverse time scales—from near real-time updates to the analysis of long-term environmental trends (Zhu et al., 2017). Airborne platforms, on the other hand, offer higher spatial resolution over more localized areas, making them ideal for detailed studies that require fine-grained spatial information.
@@ 2.1. Research Questions
Finally in Section 5, R7-8 take a holistic view of the accomplishments in this field, discussing key achievements and, in line with the current NewSpace era, the potential of AI on the edge for RS missions.
  • RQ7: Given limited FPGA resources, which ML algorithms are suitable, and which RS applications can this platform support?
  • RQ8: How are AI-powered edge computing solutions transforming RS missions?
To answer these research questions, we conduct a systematic search procedure adapted from the method proposed by the PRISMA 2020 guidelines (Page et al., 2021). The procedure starts by defining a prompt to frame for the population of records of interest. After collecting the results from the chosen research databases, we perform an initial screening pass via metadata analysis to filter out undesired records. The second screening pass involves reviewing the complete records, excluding any not clearly identified as being outside the survey’s scope based on the metadata alone. The final curated collection of studies is enriched with additional articles coming from private libraries.
@@ 2.2. Search Procedure
TOPICS: [ (Machine Learning OR Deep Learning OR Neural Network OR Convolutional Neural Network OR Recurrent Neural Network OR ML OR DL OR NN OR CNN OR RNN) AND
@@
OR Remote Sensing OR Earth Science OR LiDAR OR SAR OR UAV OR Sentinel) AND
@@
YEAR PUBLISHED: 20142014-$2024$, LANGUAGE: ENGLISH
The central part of the prompt consists of three expressions framing for any ML model, the usage of RS data in EO, and the explicit mention of FPGA devices. We choose not to consider articles published more than 10 years ago. Indeed, the FPGA technology, the state-of-the-art DL models, and even the tackled problems have evolved rapidly. Therefore, work from more than 10 years ago is often too different from modern studies to be compared fairly. After retrieving the $119$ results of the prompt, we proceed to screen the articles to filter mismatches. To this effect, we pick 4 Exclusion Criteria (EC):
@@
While EC1 and EC2 are self-explanatory, EC3 mostly rules out non-EO applications of remote sensing technology, e.g., LiDAR for autonomous driving or SAR calibration. EC4 excludes studies with no experiments or details about the FPGA implementation. For example, EC4 eliminates records mentioning FPGA hardware as a possible solution or using a different FPGA acronym, e.g., fast patch-free global learning. We design this filtering process to retain only the most relevant, experiment-based studies that reflect the current state of FPGA-based ML in Remote Sensing. After screening, $46$ articles are kept for review for a total of $48$ records when adding the records from private libraries. Each article included in the survey is thoroughly read and tagged in Zotero (Corporation for Digital Scholarship, 2024). In particular, we use Zotero’s tagging system and API to systematically report the experiments’ specificities and performance metrics. The entire database extracted via Zotero’s API, along with the code used for analysis, figures, and main taxonomy tables, is available on GitHub at https://github.com/CedricLeon/Survey_RS-ML-FPGA. Our repository also hosts supplemental material, such as extended methodological considerations and the PRISMA 2020 flow diagram.
@@ 3. Mapping the Research Landscape: Applications, Models, and Hardware
In this section, we analyze the motivations, tackled problems, and solutions presented by each surveyed study.